# Each Test Image Deserves A Specific Prompt: Continual Test-Time Adaptation for 2D Medical Image Segmentation

### v2

@@ Abstract
footnotetext: Corresponding authors: Yong Xia and Yongsheng Pan.
@@ 1 Introduction
Semantic segmentation serves as a pivotal component in medical image processing to outline specific anatomical structures. However, the segmentation model pre-trained on the source dataset may be hindered from deploying in real-world applications due to the commonly existing distribution shift, which is typically caused by the variations in imaging protocols, operators, and scanners [36, 16]. Test-time adaptation (TTA) has emerged as a promising paradigm to replace domain adaptation (DA) [50, 56, 46, 5], as TTA requires only test data during the inference phase, while DA necessitates a sufficiently large and representative target dataset [49]. The mainstream TTA methods [44, 42, 52, 51, 32, 33, 31, 39] leverage the self-supervised tasks to construct auxiliary losses, such as entropy minimization [44] and self-training [39], to update the parameters of pre-trained models. Although these TTA methods successfully alleviate the domain gap and performance degradation, they are based on the assumption that the target domains exhibit static distributions. Unfortunately, the target domains are continuously changing rather than static in most real-world scenarios. This was first presented as the continual test-time adaptation (CTTA) setup [45] that addresses TTA on a series of distribution shifts. Under this setup, pre-trained models are required to adapt to several distinct domains and therefore are more prone to suffer error accumulation and catastrophic forgetting. As mentioned above, many TTA methods rely on self-supervised tasks and may be affected by the noisy losses due to unreliable supervision, leading to error accumulation [24, 2]. Catastrophic forgetting typically occurs when the model is continually trained on new domains, resulting in unexpected performance degradation [21, 32]. Some research has been studied in addressing these two challenges based on optimizing the loss functions for better supervision [6, 45, 33, 45, 49] (see Figure 1 (a)) or prevent the model parameters from excessive alterations by model resetting or introducing regularizers [45, 33, 32, 41] (see Figure 1 (b)). Despite their efforts, loss optimization cannot ensure a robust training process to avoid catastrophic forgetting during the long-term inference, insufficient or excessive updating of the model parameters may lead to inadequate adaptation or error accumulation. Since the pre-trained model is updated by the self-supervised loss where the noise cannot be completely removed, these methods may still suffer from these two issues. Considering the above limitations in CTTA, updating the pre-trained model during the inference phase seems to be inappropriate. To this end, we propose the Visual Prompt-based Test-Time Adaptation (VPTTA) method to avoid error accumulation and catastrophic forgetting associated with updating model by freezing the model parameters and conduct adaptation by learning a specific visual prompt for each test image (see Figure 1 (c)). The prompt can adapt each test image to the frozen pre-trained model to take into account both adaptation ability and knowledge retaining. We attempt to achieve this through answering the following three questions.
@@ 2.1 Test-time Adaptation (TTA)
TTA aims to adapt the model pre-trained on the source domain to the test data during inference in a source-free and online manner [42, 44]. The mainstream TTA methods are based on model updating, which constructs self-supervised auxiliary tasks to guide the model to learn on the test data [44, 42, 52, 28, 4, 39, 31, 13]. Sun et al. [42] utilized an auxiliary task to predict the rotation to assist the model learning on the test data. Wang et al. [44] presented a test-time entropy minimization scheme to reduce generalization error by reducing the entropy of model predictions on test data. Some other methods are based on BN statistics, which alleviate the distribution shifts by modifying the statistics within BN layers [48, 29, 27]. Mirza et al. [27] constantly adapted the statistics of the BN layers to modify the feature representations of the model. These TTA methods improve the adaptation performance on the static target domain but neglect the continually changing target domains in most real-world scenarios. Recently, the continual test-time adaptation (CTTA) [45] is proposed to tackle TTA under a series of continually changing target domains. Due to long-term inference, both error accumulation [24, 2] and catastrophic forgetting [32, 21] are more prone to occur in this setup. Many methods [45, 41, 33, 6, 32, 49, 51] optimize the self-supervised losses or prevent excessive alterations to the model parameters via model resetting or regularizer to improve this situation but are still impacted by these issues. Wang et al. [45] refined pseudo labels by the weight- and augmentation-averaged predictions and randomly restored a small part of model parameters. Niu et al. [33] removed partial noisy samples with large gradients and encouraged model weights toward a flat minimum. Yang et al. [49] designed a dynamic strategy to adjust the learning rate for each test image. Zhang et al. [51] presented to adaptively fuse source and test statistics and improve the entropy minimization loss to a generalized one. In contrast, our proposed VPTTA employs a learnable prompt for each test image to mitigate distribution shifts during inference, requiring no changes to the pre-trained model and therefore alleviating the error accumulation and catastrophic forgetting.
@@
Figure 2: Overview of our VPTTA. For each test image, (1) the Fast Fourier Transform (FFT) is first applied to transform it into the frequency domain, where the low-frequency component of amplitude is used to query in a memory bank to initialize the current prompt, and then the amplitude is multiplied with the prompt at the low-frequency component and transformed back to the spatial domain using the Inverse Fast Fourier Transform (IFFT). (2) The memory bank is built on the previous low-frequency components and their corresponding prompts and is updated using the First In First Out (FIFO) strategy. (3) We convert the source statistics stored in BN layers into the warm-up statistics and calculate the absolute distance between warm-up and target statistics as the loss to fine-tune the prompt. (4) Finally, we feed the image corrected by its fine-tuned prompt to the pre-trained model to produce the output. ’Stat’: Abbreviation of ’Statistics’.
@@ 3.1 Problem Definition and Method Overview
Let the labeled source domain dataset and unlabeled testing target dataset be $\mathbb{D}^{s}=\left\{X_{i}^{s},Y_{i}^{s}\right\}_{i=1}^{N^{s}}$ and $\mathbb{D}^{t}=\left\{X_{i}^{t}\right\}_{i=1}^{N^{t}}$, respectively, where $X_{i}^{*}\in\mathbb{R}^{H\times W\times C}$ is the $i$-th image with C channels and size of $H\times W$, and $Y_{i}^{*}$ is its label. Our goal is to learn a low-frequency prompt $\mathcal{P}_{i}$ for each test image $X_{i}^{t}$ to obtain the modified image $\tilde{X}_{i}^{t}$ that can be adapted to the model $f_{\theta}:X\rightarrow Y$ pre-trained on $\mathbb{D}^{s}$. To accomplish this goal, our VPTTA considers three key facets: prompt design, prompt initialization, and prompt training. For prompt design, the prompt is designed to change the low-frequency component of the amplitude of the test image. For prompt initialization, we adopt a memory bank to initialize the current prompt with previous ones. For prompt training, we employ the absolute distance to train the prompt to align the statistics between features extracted from test images and the batch normalization (BN) layers and also present a statistics-fusion-based warm-up mechanism. An overview of our VPTTA is illustrated in Figure 2. We now delve into its details.
@@ 3.2 Prompt Design: Low-frequency Prompt
where $\odot$ is the element-wise multiplication, and $OnePad$ means padding one around $\mathcal{P}_{i}$ to expand it to the size of $H\times W$. Under the control of the small coefficient $\alpha\in(0,1)$, we can ensure that the prompt $\mathcal{P}_{i}$ remains lightweight and focuses on the low-frequency component. Simultaneously, we recognize that the low-rank decomposition is widely employed in the field of data compression [22] and model fine-tuning [18]. Consequently, we also design a low-rank prompt $\mathcal{P}_{i}^{LoRa}=\mathcal{B}_{i}@\mathcal{A}_{i}$ as a variant (denoted as VPTTA-LR), where $\mathcal{B}_{i}\in\mathbb{R}^{H\times r\times C}(r\ll W)$, $\mathcal{A}_{i}\in\mathbb{R}^{r\times W\times C}(r\ll H)$, and @@@@ denotes the matrix multiplication operation. In this way, the adapted image can be formalized as $\tilde{X}_{i}^{t}=X_{i}^{t}+\mathcal{P}_{i}^{LoRa}$.
@@ 3.3 Prompt Initialization: Memory Bank
10:  Forward $f_{\theta}(\tilde{X}_{i}^{t})$ and calculate the warm-up statistics $\mu_{w}$ and $\sigma_{w}$ by Eq. (4) and (5) to replace the $\mu_{s}$ and $\sigma_{s}$ stored in $f_{\theta}$
@@ 3.4 Prompt Training: Statistics Alignment
where $j$ indicates the $j$-th BN layer in $f_{\theta}$. However, aligning the source and target domains within a limited number of iterations (one in this study) provides challenges due to the empty memory bank and undiminished distribution shifts at the beginning of the inference phase. Therefore, we design a warm-up mechanism to address this issue by simulating the warm-up statistics (i.e., $\mu_{w}$ and ${\sigma}_{w}$) as follows:
$\mu_{w}=\lambda*\mu_{t}+(1-\lambda)*\mu_{s},$ (4)
$\sigma_{w}=\lambda*\sigma_{t}+(1-\lambda)*\sigma_{s},$ (5)
@@ 4.1 Datasets and Evaluation Metrics
The OD/OC segmentation dataset comprises five public datasets collected from different medical centres, denoted as domain A (RIM-ONE-r3 [11]), B (REFUGE [34]), C (ORIGA [53]), D (REFUGE-Validation/Test [34]), and E (Drishti-GS [40]). There are 159, 400, 650, 800, and 101 images from these datasets. We cropped a region of interest (ROI) centering at OD with size of $800\times 800$ for each image following [19], and each ROI is further resized to $512\times 512$ and normalized by min-max normalization. The Dice score metric ($DSC$) is utilized for evaluation in this task. The polyp segmentation dataset consists of four public datasets collected from different medical centres, denoted as domain A (BKAI-IGH-NeoPolyp [30]), B (CVC-ClinicDB [1]), C (ETIS-LaribPolypDB [38]), and D (Kvasir-Seg [20]), which have 1000, 612, 196, and 1000 images, respectively. We followed [9] to resize each image to $352\times 352$ and normalize the resized image by the statistics computed on ImageNet. The $DSC$, enhanced-alignment metric ($E^{max}_{\phi}$[8], and structural similarity metric ($S_{\alpha}$[7] are utilized for evaluation in this task.
@@ 4.3 Experimental Results
We compare our VPTTA with the ’Source Only’ baseline (training on the source domain and testing without adaptation), and six competing methods, including a pseudo-label-based method (CoTTA [45]), two entropy-based methods (TENT-continual [44] and SAR [33]), a method dynamically adjusting the learning rate (DLTTA [49]), a method combining entropy-loss and BN statistics fusion (DomainAdaptor [51]), and a BN statistics-based method (DUA [27]) on two segmentation tasks. Specifically, we re-implemented all the competing methods using the same baseline as our VPTTA and combined DLTTA with the entropy loss proposed in TENT [49]. The result of each domain is calculated by using the current domain as the source domain and applying the pre-trained model on the left domains. Comparison on the OD/OC segmentation task. The results of the proposed VPTTA, the “Source Only” baseline, and six competing methods on the OD/OC segmentation task are detailed in Table 1. Obviously, all competing methods outperform the “Source Only” baseline, indicating the effectiveness of adaptation towards the distribution shifts. It can be found that our VPTTA achieves the best overall performance across all domains, underscoring its superior applicability and robustness. Meanwhile, our VPTTA also shows competitive performance in domains where most methods exhibit satisfactory performance (such as D) and also performs well in domains where other methods exhibit suboptimal results (such as C).
@@
Figure 3: Visualization of the original images, estimated prompts, and adapted images on the OD/OC segmentation task. We normalize the prompts to [0, 1] for better visualization. The DSC of applying the frozen source model on the original and adapted images is displayed below each image. We also show an example of each source domain on the left side of this diagram. ’Ori’: Abbreviation of ’Original’.
@@
Comparison on the polyp segmentation task. We conducted the comparison experiment on the polyp segmentation task under the same experimental setup, as shown in Table 2. In contrast to the results of the OD/OC segmentation task, the methods based on model updating (i.e., TENT-continual, CoTTA, DLTTA, SAR, and DomainAdaptor) exhibit significantly worse performance in this task and are even inferior to the ’Source Only’ baseline. It can be attributed to the fact that the polyps are relatively hidden, and the models tend to segment almost nothing rather than poorly segment under the impact of distribution shifts, resulting in confident and low entropy predictions but wrong gradients for these methods. Since DUA is backward-free, it still performs well in this task. However, DUA also fails when using domain A as the source domain and has limited performance due to the absence of training for the current test image. Our VPTTA simultaneously benefits from training on the current test image and refraining from updating the model, thus surpassing all competing methods and demonstrating advanced performance. Visualization of prompts and adapted images. We visualized the estimated prompts and adapted images on the OD/OC segmentation task in Figure 3 and displayed the DSC values predicted on the original and adapted images below the corresponding images. It can be observed that: (1) Applying the prompts to the original test images can achieve significant performance gain; (2) Images from different target domains have similar prompts, proving that involving the memory bank to initialize the current prompt using previous ones is reasonable; (3) The appearances of adapted images are close to the source domain, explicitly demonstrating the reduction of distribution shifts via our VPTTA.
@@
Table 4: Performance of our VPTTA and four competing methods on the OD/OC segmentation task in a long-term continual test-time adaptation. The red numbers indicate the performance gain relative to the ’Source Only’ baseline. The performance degradation is calculated between the overall average DSC and the average DSC of round 1. ’Ave’: Abbreviation of ’Average’.
@@
Comparison in a long-term continual test-time adaptation. To evaluate our VPTTA in a long-term continual test-time adaptation [45], we conducted the experiments for the OD/OC segmentation task multiple rounds and reported the results in Table 4. Since DUA is a backward-free method and DomainAdaptor resets the pre-trained model before each adaptation, they are unsuitable to be evaluated in this experiment. We compared our VPTTA with TENT, CoTTA, DLTTA, and SAR, all of which involve continuous training of the pre-trained model. The performance across different rounds with the same source domain reflects the extent of catastrophic forgetting in the model, and the average performance across multiple rounds highlights the error accumulation of the model within a series of distinct domains. The results reveal that both optimizing loss functions (CoTTA, DLTTA, SAR) and resetting the model (CoTTA, SAR) are effective for mitigating error accumulation and catastrophic forgetting, while TENT with vanilla entropy-minimization loss suffers from severe performance degradation. Additionally, after multiple rounds of testing, our VPTTA consistently maintains superior performance with minimal performance degradation, emphasizing the effectiveness of training the prompt over updating the pre-trained model. Ablation study. To evaluate the contributions of the low-frequency prompt, memory-bank-based initialization, and warm-up mechanism, we conducted ablation studies on the OD/OC segmentation task, as shown in Table 3. Note that Eq. (3)/(6) is used as the loss function when training the prompt without/with our warm-up mechanism. The results show that (1) using the low-frequency prompt only has limited performance gain due to the poorly trained prompt; (2) both the memory bank and warm-up mechanism are beneficial to training the prompt; (3) the best performance is achieved when the memory bank and warm-up mechanism are jointly used, which means they are complementary.
@@
Analysis of hyperparameter $\bm{\alpha}$, $\bm{S}$, $\bm{K}$, and $\bm{\tau}$. We further conducted the experiments for the OD/OC segmentation task to discuss the hyper-parameters used in our VPTTA and displayed the results in Figure 4 and Figure 5. It can be seen that $\alpha=0.01$ is optimal, as an excessively small prompt makes adaptation challenging, while an overly large prompt hinders effective training. Similar trends are also reflected in $S$ and $K$: as they increase, the DSC values improve, while the excessively large $S$ and $K$ may introduce unrelated samples, consequently affecting the initialization. With the increase of $\tau$, performance gradually approaches the upper bound and tends to remain stable, suggesting that the warm-up has reached saturation. Therefore, we finally chose $\alpha=0.01$, $S=40$, $K=16$, and $\tau=5$ as the best configuration.

### v4

@@ Abstract
Yong Xia and Yongsheng Pan are the corresponding authors. This work was supported in part by the National Natural Science Foundation of China under Grants 62171377, in part by STI2030-Major Project under Grants 2022ZD0213100, in part by Shenzhen Science and Technology Program under Grants JCYJ20220530161616036, and in part by the Ningbo Clinical Research Center for Medical Imaging under Grant 2021L003 (Open Project 2022LYKFZD06).
@@ 1 Introduction
Semantic segmentation serves a pivotal role in medical image processing to outline specific anatomical structures. However, the segmentation model pre-trained on the source dataset may be hindered from deploying in real-world applications due to the commonly existing distribution shift, which is typically caused by the variations in imaging protocols, operators, and scanners [36, 16]. Test-time adaptation (TTA) has emerged as a promising paradigm to replace domain adaptation (DA) [50, 57, 46, 5], as TTA requires only test data during the inference phase, while DA necessitates a sufficiently large and representative target dataset [49]. The mainstream TTA methods [44, 42, 52, 51, 32, 33, 31, 39] leverage the self-supervised tasks to construct auxiliary losses, such as entropy minimization [44] and self-training [39], to update the parameters of pre-trained models. Although these TTA methods successfully alleviate the domain gap and performance degradation, they are based on the assumption that the target domains exhibit static distributions. Unfortunately, the target domains are continuously changing rather than static in most real-world scenarios.
This was first presented as the continual test-time adaptation (CTTA) setup [45] that addresses TTA on a series of distribution shifts. Under this setup, pre-trained models are required to adapt to several distinct domains and therefore are more prone to suffer error accumulation and catastrophic forgetting. As mentioned above, many TTA methods rely on self-supervised tasks and may be affected by the noisy losses due to unreliable supervision, leading to error accumulation [24, 2]. Catastrophic forgetting typically occurs when the model is continually trained on new domains, resulting in unexpected performance degradation [21, 32]. Some research has been studied in addressing these two challenges based on optimizing the loss functions for better supervision [6, 45, 33, 45, 49] (see Figure 1 (a)) or prevent the model parameters from excessive alterations by model resetting or introducing regularizers [45, 33, 32, 41] (see Figure 1 (b)). Despite their efforts, loss optimization cannot ensure a robust training process to avoid catastrophic forgetting during the long-term inference, insufficient or excessive updating of the model parameters may lead to inadequate adaptation or error accumulation. Since the pre-trained model is updated by the self-supervised loss where the noise cannot be completely removed, these methods may still suffer from these two issues.
Considering the above limitations in CTTA, updating the pre-trained model during inference seems to be inappropriate. To this end, we propose the Visual Prompt-based Test-Time Adaptation (VPTTA) method to avoid error accumulation and catastrophic forgetting associated with updating model by freezing the model parameters and conduct adaptation by learning a specific visual prompt for each test image (see Figure 1 (c)). The prompt can adapt each test image to the frozen pre-trained model to take into account both adaptation ability and knowledge retaining. We attempt to achieve this by answering the following three questions.
@@ 2.1 Test-time Adaptation (TTA)
TTA aims to adapt the model pre-trained on the source domain to the test data during inference in a source-free and online manner [42, 44]. The mainstream TTA methods are based on model updating, which constructs self-supervised auxiliary tasks to guide the model to learn on the test data [44, 42, 52, 28, 4, 39, 31, 13, 53]. Sun et al. [42] utilized an auxiliary task to predict the rotation to assist the model learning on the test data. Wang et al. [44] presented a test-time entropy minimization scheme to reduce generalization error by reducing the entropy of model predictions on test data. Zhang et al. [53] introduced a supplementary network to fit the test data, which can work in a plug-and-play fashion, necessitating minimal hyperparameter tuning. Some other methods are based on BN statistics, which alleviate the distribution shifts by modifying the statistics within BN layers [48, 29, 27]. Mirza et al. [27] constantly adapted the statistics of the BN layers to modify the feature representations of the model. These TTA methods improve the adaptation performance on the static target domain but neglect the continually changing target domains in most real-world scenarios.
Recently, the continual test-time adaptation (CTTA) [45] is proposed to tackle TTA under a series of continually changing target domains. Due to long-term inference, both error accumulation [24, 2] and catastrophic forgetting [32, 21] are more prone to occur in this setup. Many methods [45, 41, 33, 6, 32, 49, 51] optimize the self-supervised losses or prevent excessive alterations to the model parameters via model resetting or regularizer to improve this situation but are still impacted by these issues. Wang et al. [45] refined pseudo labels by the weight- and augmentation-averaged predictions and randomly restored a small part of model parameters. Niu et al. [33] removed partial noisy samples with large gradients and encouraged model weights toward a flat minimum. Yang et al. [49] designed a dynamic strategy to adjust the learning rate for each test image. Zhang et al. [51] presented to adaptively fuse source and test statistics and improve the entropy minimization loss to a generalized one. In contrast, our proposed VPTTA employs a learnable prompt for each test image to mitigate distribution shifts during inference, requiring no changes to the pre-trained model and therefore alleviating the error accumulation and catastrophic forgetting.
@@
Figure 2: Overview of our VPTTA. For each test image, (1) the Fast Fourier Transform (FFT) is first applied to transform it into the frequency domain, where the low-frequency component of amplitude is used to query in a memory bank to initialize the current prompt, and then the amplitude is multiplied with the prompt at the low-frequency component and transformed back to the spatial domain using the Inverse Fast Fourier Transform (IFFT). (2) The memory bank is built on the previous low-frequency components and their corresponding prompts and is updated using the First In First Out (FIFO) strategy. (3) We convert the source statistics stored in BN layers into the warm-up statistics and calculate the absolute distance between warm-up and target statistics as the loss to fine-tune the prompt. (4) Finally, we feed the image corrected by its fine-tuned prompt to the pre-trained model to produce the output. ‘Stat’: Abbreviation of ‘Statistics’.
@@ 3.1 Problem Definition and Method Overview
Let the labeled source domain dataset and unlabeled testing target dataset be $\mathbb{D}^{s}=\left\{X_{i}^{s},Y_{i}^{s}\right\}_{i=1}^{N^{s}}$ and $\mathbb{D}^{t}=\left\{X_{i}^{t}\right\}_{i=1}^{N^{t}}$, respectively, where $X_{i}^{*}\in\mathbb{R}^{H\times W\times C}$ is the $i$-th image with C channels and size of $H\times W$, and $Y_{i}^{*}$ is its label. Our goal is to learn a low-frequency prompt $\mathcal{P}_{i}$ for each test image $X_{i}^{t}$ to obtain the modified image $\tilde{X}_{i}^{t}$ that can be adapted to the model $f_{\theta}:X\rightarrow Y$ pre-trained on $\mathbb{D}^{s}$.
To accomplish this goal, our VPTTA considers three key facets: prompt design, prompt initialization, and prompt training. For prompt design, the prompt is designed to change the low-frequency component of the amplitude of the test image. For prompt initialization, we adopt a memory bank to initialize the current prompt with previous ones. For prompt training, we employ the absolute distance to train the prompt to align the statistics between features extracted from test images and the batch normalization (BN) layers and also present a statistics-fusion-based warm-up mechanism. An overview of our VPTTA is illustrated in Figure 2. We now delve into its details.
@@ 3.2 Prompt Design: Low-frequency Prompt
where $\odot$ is the element-wise multiplication, and $OnePad$ means padding one around $\mathcal{P}_{i}$ to expand it to the size of $H\times W$. Under the control of the small coefficient $\alpha\in(0,1)$, we can ensure that the prompt $\mathcal{P}_{i}$ remains lightweight and focuses on the low-frequency component.
Simultaneously, we recognize that the low-rank decomposition is widely employed in the field of data compression [22] and model fine-tuning [18]. Consequently, we also design a low-rank prompt $\mathcal{P}_{i}^{LoRa}=\mathcal{B}_{i}@\mathcal{A}_{i}$ as a variant (denoted as VPTTA-LR), where $\mathcal{B}_{i}\in\mathbb{R}^{H\times r\times C}(r\ll W)$, $\mathcal{A}_{i}\in\mathbb{R}^{r\times W\times C}(r\ll H)$, and @@@@ denotes the matrix multiplication operation. In this way, the adapted image can be formalized as $\tilde{X}_{i}^{t}=X_{i}^{t}+\mathcal{P}_{i}^{LoRa}$.
@@ 3.3 Prompt Initialization: Memory Bank
10:  Forward $f_{\theta}(\tilde{X}_{i}^{t})$ and calculate the warm-up statistics $\mu_{w}$ and $\sigma_{w}$ by Eq. (4) to replace the $\mu_{s}$ and $\sigma_{s}$ stored in $f_{\theta}$
@@ 3.4 Prompt Training: Statistics Alignment
where $j$ indicates the $j$-th BN layer in $f_{\theta}$.
However, aligning the source and target domains within a limited number of iterations (one in this study) provides challenges due to the empty memory bank and undiminished distribution shifts at the beginning of the inference phase. Therefore, we design a warm-up mechanism to address this issue by simulating the warm-up statistics (i.e., $\mu_{w}$ and ${\sigma}_{w}$) as follows:
$\mu_{w}=\lambda\mu_{t}+(1-\lambda)\mu_{s},\,\sigma_{w}=\lambda\sigma_{t}+(1-\lambda)\sigma_{s},$ (4)
@@
Table 1: Performance of our VPTTA, ‘Source Only’ baseline, and six competing methods on the OD/OC segmentation task. The best and second-best results in each column are highlighted in bold and underline, respectively.
@@
Table 2: Performance of our VPTTA, ‘Source Only’ baseline, and six competing methods on the polyp segmentation task. The best and second-best results in each column are highlighted in bold and underline, respectively.
@@ 4.1 Datasets and Evaluation Metrics
The OD/OC segmentation dataset comprises five public datasets collected from different medical centres, denoted as domain A (RIM-ONE-r3 [11]), B (REFUGE [34]), C (ORIGA [54]), D (REFUGE-Validation/Test [34]), and E (Drishti-GS [40]). There are 159, 400, 650, 800, and 101 images from these datasets. We cropped a region of interest (ROI) centering at OD with size of $800\times 800$ for each image following [19], and each ROI is further resized to $512\times 512$ and normalized by min-max normalization. The Dice score metric ($DSC$) is utilized for evaluation in this task.
The polyp segmentation dataset consists of four public datasets collected from different medical centres, denoted as domain A (BKAI-IGH-NeoPolyp [30]), B (CVC-ClinicDB [1]), C (ETIS-LaribPolypDB [38]), and D (Kvasir-Seg [20]), which have 1000, 612, 196, and 1000 images, respectively. We followed [9] to resize each image to $352\times 352$ and normalize the resized image by the statistics computed on ImageNet. The $DSC$, enhanced-alignment metric ($E^{max}_{\phi}$[8], and structural similarity metric ($S_{\alpha}$[7] are utilized for evaluation in this task.
@@ 4.3 Experimental Results
We compare our VPTTA with the ‘Source Only’ baseline (training on the source domain and testing without adaptation), and six competing methods, including a pseudo-label-based method (CoTTA [45]), two entropy-based methods (TENT-continual [44] and SAR [33]), a method dynamically adjusting the learning rate (DLTTA [49]), a method combining entropy-loss and BN statistics fusion (DomainAdaptor [51]), and a BN statistics-based method (DUA [27]) on two segmentation tasks. Specifically, we re-implemented all the competing methods using the same baseline as VPTTA and combined DLTTA with the entropy loss proposed in TENT [49]. The result of each domain is calculated by using the current domain as the source domain and applying the pre-trained model on the left domains.
Comparison on the OD/OC segmentation task. The results of the proposed VPTTA, the “Source Only” baseline, and six competing methods on the OD/OC segmentation task are detailed in Table 1. Obviously, all competing methods outperform the “Source Only” baseline, indicating the effectiveness of adaptation towards the distribution shifts. It can be found that our VPTTA achieves the best overall performance across all domains, underscoring its superior applicability and robustness. Meanwhile, our VPTTA also shows competitive performance in domains where most methods exhibit satisfactory performance (such as D) and also performs well in domains where other methods exhibit suboptimal results (such as C).
@@
Figure 3: Visualization of the original images, estimated prompts, and adapted images on the OD/OC segmentation task. We normalize the prompts to [0, 1] for better visualization. The DSC of applying the frozen source model on the original and adapted images is displayed below each image. We also show an example of each source domain on the left side of this diagram. ‘Ori’: Abbreviation of ‘Original’.
@@
Comparison on the polyp segmentation task. We conducted the comparison experiment on the polyp segmentation task under the same experimental setup, as shown in Table 2. In contrast to the results of the OD/OC segmentation task, the methods based on model updating (i.e., TENT-continual, CoTTA, DLTTA, SAR, and DomainAdaptor) exhibit significantly worse performance in this task and are even inferior to the ‘Source Only’ baseline. It can be attributed to the fact that the polyps are relatively hidden, and the models tend to segment almost nothing rather than poorly segment under the impact of distribution shifts, resulting in confident and low entropy predictions but wrong gradients for these methods. Since DUA is backward-free, it still performs well in this task. However, DUA also fails when using domain A as the source domain and has limited performance due to the absence of training for the current test image. Our VPTTA simultaneously benefits from training on the current test image and refraining from updating the model, thus surpassing all competing methods and demonstrating advanced performance.
Visualization of prompts and adapted images. We visualized the estimated prompts and adapted images on the OD/OC segmentation task in Figure 3 and displayed the DSC values predicted on the original and adapted images below the corresponding images. It can be observed that: (1) Applying prompts to the original test images can achieve significant performance gain; (2) Images from different target domains have similar prompts, proving that involving the memory bank to initialize the current prompt using previous ones is reasonable; (3) The appearances of adapted images are close to the source domain, explicitly demonstrating the reduction of distribution shifts via our VPTTA.
@@
Table 4: Performance of our VPTTA and four competing methods on the OD/OC segmentation task in a long-term continual test-time adaptation. The red numbers indicate the performance gain relative to the ‘Source Only’ baseline. The performance degradation is calculated between the overall average DSC and the average DSC of round 1. ‘Ave’: Abbreviation of ‘Average’.
@@
Comparison in a long-term continual test-time adaptation. To evaluate our VPTTA in a long-term continual test-time adaptation [45], we conducted the experiments for the OD/OC segmentation task multiple rounds and reported the results in Table 4. Since DUA is a backward-free method and DomainAdaptor resets the pre-trained model before each adaptation, they are unsuitable to be evaluated in this experiment. We compared our VPTTA with TENT, CoTTA, DLTTA, and SAR, all of which involve continuous training of the pre-trained model. The performance across different rounds with the same source domain reflects the extent of catastrophic forgetting in the model, and the average performance across multiple rounds highlights the error accumulation of the model within a series of distinct domains. The results reveal that both optimizing loss functions (CoTTA, DLTTA, SAR) and resetting the model (CoTTA, SAR) are effective for mitigating error accumulation and catastrophic forgetting, while TENT with vanilla entropy-minimization loss suffers from severe performance degradation. Additionally, after multiple rounds of testing, our VPTTA maintains superior performance with minimal performance degradation, emphasizing the effectiveness of training prompts over updating the pre-trained model.
Ablation study. To evaluate the contributions of the low-frequency prompt, memory-bank-based initialization, and warm-up mechanism, we conducted ablation studies on the OD/OC segmentation task, as shown in Table 3. Note that Eq. (3)/(5) is used as the loss function when training the prompt without/with our warm-up mechanism. The results show that (1) using the low-frequency prompt only has limited performance gain due to the poorly trained prompt; (2) both the memory bank and warm-up mechanism are beneficial to training the prompt; (3) the best performance is achieved when the memory bank and warm-up mechanism are jointly used, which means they are complementary.
@@
Analysis of hyperparameter $\bm{\alpha}$, $\bm{S}$, $\bm{K}$, and $\bm{\tau}$. We conducted the experiments for the OD/OC segmentation task to discuss the hyper-parameters used in our VPTTA and displayed the results in Figure 4 and Figure 5. It can be seen that $\alpha=0.01$ is optimal, as an excessively small prompt makes adaptation challenging, while an overly large prompt hinders effective training. Similar trends are also reflected in $S$ and $K$: as they increase, the DSC values improve, while the excessively large $S$ and $K$ may introduce unrelated samples, affecting the initialization. With the increase of $\tau$, performance gradually approaches the upper bound and tends to remain stable, suggesting that the warm-up has reached saturation. Therefore, we finally chose $\alpha=0.01$, $S=40$, $K=16$, and $\tau=5$ as the best configuration.